# A Comprehensive Characterization of Galaxy-Cool CGM Connections at $z<0.4$ with DESI Year 1 Data

### v1

@@ Abstract
We investigate the relationships between the properties of the cool circumgalactic medium (CGM), traced by Ca II absorption lines, and those of galaxies at $z<0.4$ by utilizing a galaxy-quasar pair sample compiled from the Year 1 data of the Dark Energy Spectroscopic Instrument (DESI). This large dataset, containing $\sim 900,000$ galaxy-quasar pairs within $200\,\rm kpc$, enables us to obtain composite spectra with sensitivity reaching to $\rm m\AA$ level and to explore the Ca II absorption as a function of stellar mass, SFR, redshift, and galaxy types, including AGNs. Our results show a positive correlation between the absorption strength and stellar mass of star-forming galaxies with $\langle W_{0}^{\rm Ca\ II}\rangle\propto M_{*}^{0.5}$ over three orders of magnitude in stellar mass from $\sim 10^{8}$ to $10^{11}\,M_{\odot}$, while such a mass dependence is weaker for quiescent galaxies. For galaxies with similar mass, we find that Ca II absorption is stronger around star-forming galaxies than around quiescent galaxies especially within the inner regions ($<30\,\rm kpc$) of the halos. Among star-forming galaxies, the Ca II absorption further correlates with SFR, following $\propto\mathrm{SFR^{0.3}}$. However, in contrast to the results at higher redshifts, we find that stronger absorption around star-forming galaxies is not preferentially observed along the minor axis of galaxies, indicating a possible redshift evolution of CGM dynamics resulting from galactic feedback. Moreover, no significant difference between the properties of the cool gas around AGNs and galaxies is detected. Finally, we measure the absorption profiles with respect to the virial radius of dark matter halos and estimate the total Ca II mass in the CGM. The results show that the CGM contains a metal mass comparable to the metal mass in the ISM of galaxies.
@@ 1 Introduction
The circumgalactic medium (CGM) contains signatures of complex physical processes that regulate gas flowing in and out of galaxies and drive galaxy evolution (see Tumlinson et al., 2017; Péroux & Howk, 2020; Faucher-Giguère & Oh, 2023, for reviews on the topic of the CGM). Therefore, observed properties of the CGM and its connection to the properties of galaxies have been considered as crucial constraints on the models of galaxy evolution, especially the subgrid models implemented for the feedback processes driven by the explosions of massive stars and the activities of supermassive black holes (e.g., Liang et al., 2016; Ford et al., 2016; Anglés-Alcázar et al., 2017; Oppenheimer et al., 2018). To probe the properties of the CGM, absorption line spectroscopy has been widely used (e.g., Bahcall & Spitzer, 1969; Bergeron, 1986; Chen et al., 2010; Nielsen et al., 2013; Tumlinson et al., 2013; Zhu et al., 2014; Lan et al., 2014; Anand et al., 2021; Bouché et al., 2025). By observing the spectra of background sources intercepting the CGM of galaxies, one can detect absorption line features induced by various transitions of elements, and investigate the properties of the circumgalactic gas, such as the radial distribution, kinematics, abundance, metallicity, and gas density, and explore how they correlate with the properties of the galaxies (e.g., Zych et al., 2009; Zhu et al., 2014; Lan et al., 2014; Nielsen et al., 2017; Pointon et al., 2019; Lan, 2020; Anand et al., 2021; Nateghi et al., 2024; Lehner et al., 2024; Cherrey et al., 2025).
@@
Multi-wavelength observations are required to investigate the galaxy-CGM connections at different redshifts owing to the accessibility of strong absorption-line transitions. At $z<0.3$, UV spectrographs, COS and STIS (Green et al., 2012; Woodgate et al., 1998), on the Hubble Space Telescope have been used to detect absorption species, such as H I, O I, O VI, C II, C IV, Si II, and Si III, tracing gas in various physical conditions (e.g., Tumlinson et al., 2013; Werk et al., 2013, 2014; Bordoloi et al., 2014; Liang & Chen, 2014; Borthakur et al., 2015, 2016; Werk et al., 2016; Prochaska et al., 2017; Chen et al., 2020, 2020; Tchernyshyov et al., 2023; Zheng et al., 2024). These measurements have led to new explorations of the nature and origins of the multiphase CGM. However, given the sensitivity of the current instruments, only bright quasars can yield UV spectra with sufficient sensitivity for absorption line studies (e.g., Tumlinson et al., 2013). This in turn limits the number of galaxy-quasar pairs and the parameter space of galaxy properties available for galaxy-CGM correlation analysis. On the other hand, at $z\gtrsim 0.4$, optical spectra can be used to investigate the connection between galaxies and the cool CGM traced by Mg II absorption lines, one of the key accessible absorption species at optical wavelengths (e.g., Nestor et al., 2005; Chen et al., 2010; Nielsen et al., 2013; Zhu & Ménard, 2013a; Anand et al., 2021; Huang et al., 2021; Chen et al., 2023; Guha et al., 2024; Zou et al., 2024; Bouché et al., 2025). With large optical spectroscopic datasets provided by sky surveys, such as the Sloan Digital Sky Surveys (SDSS; York et al. 2000) and the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2013), statistical samples, consisting of thousands of galaxy-quasar pairs, have been built and used to measure the properties of the cool CGM as a function of galaxy properties, including stellar mass, star-formation rate (SFR) and AGN activity, across redshifts (e.g., Zhu & Ménard, 2013a; Lan et al., 2014; Lan, 2020; Anand et al., 2021; Chang et al., 2024; Wu et al., 2024), revealing correlations between the cool CGM and galaxy properties. However, observational measurements of these correlations at low redshifts are still limited, and therefore whether the relationships between galaxies and the cool CGM detected at $z\gtrsim 0.4$ exist at lower redshifts remains to be addressed.
@@
In this work, we utilize the large spectroscopic data provided by the DESI survey (DESI Collaboration et al., 2016a, b) to measure the properties of the cool CGM traced by Ca II as a function of galaxy properties, including SFR, stellar mass, redshift, and azimuthal angle, covering a parameter space that has rarely been probed previously. We describe our sample and the analysis procedure to obtain composite spectra in Section 2. In Section 3, we present our results. In Section 4, we compare our results with previous studies and discuss the implications of our findings. Finally, we conclude our results in Section 5. Throughout this paper, we adopt a flat $\Lambda$CDM model with $h=0.7$ and $\Omega_{m}=0.3$.
@@ 2.1 DESI galaxy and quasar catalogs
The Dark Energy Spectroscopic Instrument (DESI) project is designed to measure the expansion rate of the Universe across cosmic time by using galaxies, quasars, and gas as tracers of large-scale structure at different redshifts (Levi et al., 2013; DESI Collaboration et al., 2022). To this end, a dedicated spectroscopic observational program has been developed utilizing an instrument consisting of 10 spectrographs and 5000 fibers (Silber et al., 2023; Miller et al., 2024). DESI primarily obtains spectra of four types of extragalactic sources at different redshift ranges, including galaxies that can be observed during the bright time targeted by Bright Galaxy Survey ($z<0.6$; Hahn et al. 2023), luminous red galaxies ($1.00.4<z<$ 1.0; Zhou et al. 2023), emission-line galaxies ($1.60.6<z<$ 1.6; Raichoor et al. 2023) and quasars ($z<3$; Chaussidon et al. 2023). These sources are selected based on the images of the DESI Legacy Imaging Surveys (Dey et al., 2019). The wavelength coverage of DESI spectra is from 3600 to 9800 Å with the spectral resolution varying from $R\sim 2000$ to 5000 for the blue camera to the NIR camera. The raw spectroscopic data are processed with a pipeline developed by Guy et al. (2023). In addition, an automatic pipeline, called redrock [^1] https://github.com/desihub/redrock, is developed and used for determining the redshifts of the sources from the spectra (Bailey et al. 2024; see also Anand et al. 2024). In order to maximize the efficiency of the survey, dedicated algorithms and codes have also been developed for target selection (Myers et al., 2023), fiber assignments (Raichoor et al., 2024) and survey operations (Schlafly et al., 2023). Prior to the main spectroscopy survey, DESI has undergone the “Survey Validation (SV)” (DESI Collaboration et al., 2024a) to validate the performance of the target selections. The five-month spectroscopic data from Survey Validation was released in June 2023 as part of Early Data Release (DESI Collaboration et al., 2024b).
@@
Figure 1: SFR-mass distribution of the foreground galaxies. The solid line divides the samples into star-forming and quiescent galaxies. The contours mark the levels of numbers at 10, 100, 500, 1000, 5000, 8000.
@@
In this work, we use the Year 1 dataset of the DESI survey (DESI Collaboration et al., 2024), which has yielded new measurements for constraining the expansion history of the Universe from $z=0$ to 2 (DESI Collaboration et al., 2024c, d, e, f, g, h; DESI Collaboration et al., 2025). More specifically, we construct our foreground galaxy and background source samples with the following two components:
@@
  • Bright Galaxy Survey (BGS): We use galaxies with $0.40.02<z_{\mathrm{BGS}}<$ 0.4 observed in the bright galaxy survey as our foreground galaxies. The BGS sample consists of two selections, BGS bright and BGS faint samples. The BGS bright sample is selected with $r<19.5$ mag, and the BGS faint sample is selected with magnitude and color cuts (Hahn et al., 2023). Moreover, following the criteria from Hahn et al. (2023) and Lan et al. (2023), we select galaxies with $\Delta\chi^{2}\geq 40$ (the difference of $\chi^{2}$ between the second and the first best-fit spectral models from redrock).
@@
    Besides the redshift information, we use the physical properties of the BGS galaxies, including the stellar mass ($M_{*}$) and star-formation rate (SFR), provided by Siudek et al. (2024). The authors use the CIGALE package (Boquien et al., 2019) for spectral energy distribution fits to observed magnitudes in the $g$, $r$, $z$, WISE 1 and WISE 2 bands extracted by the Tractor algorithm (Lang et al., 2016). The final fitting results exclude sources with low signal-to-noise ratio in photometry and poor fitting quality. For details, we refer the readers to Siudek et al. (2024).
@@
    As shown in the figure, the BGS sample includes star-forming galaxies with $M_{*}$ down to $10^{7}-10^{8}\,M_{\odot}$ allowing us to probe the CGM of galaxies with mass across 3-4 orders of magnitude. The median mass uncertainty of the sample is 0.12 dex. For SFR, the typical uncertainties are 0.3 dex for star-forming galaxies and 8.1 dex for quiescent galaxies. We note that the reported uncertainty is in logarithmic scale. For quiescent galaxies with low estimated SFR ($\log_{10}\mathrm{SFR}/M_{\odot}\rm yr^{-1}\sim-7$), the corresponding uncertainty in linear scale is $\sim 10^{-6}M_{\odot}/\rm yr$. This precision is sufficient for classifying galaxies into star-forming and quiescent ones.
    In addition, we construct an AGN sample by selecting sources located in the AGN and composite regions (Kewley et al., 2006) of the BPT diagram (Baldwin et al., 1981). For this selection, we require a signal-to-noise ratio greater than 2 for the $\rm H\alpha$, $\rm H\beta$, [O III] 5007, and [N II] 6584 emission lines. The flux measurements for these lines are provided by the FastSpecFit [^2] https://fastspecfit.readthedocs.io/en/latest/ Spectral Synthesis and Emission-Line Catalog (FastSpecFit 2.1) (Moustakas et al., 2023).
  • Quasars: For the background sources, we use the quasar sample which is selected with a random forest algorithm combining magnitude and color information (Chaussidon et al., 2023). Similar to the BGS sample selection, we apply $\Delta\chi^{2}\geq 20$ to obtain sources with robust redrock redshift measurements (Alexander et al., 2023).
With the galaxy and quasar samples, we construct the galaxy-quasar pair sample by requiring (1) $z_{\mathrm{QSO}}-z_{\mathrm{BGS}}>0.1$ and (2) $\left(z_{\mathrm{BGS}}+1\right)\times 3800\,\mathrm{\AA}>\left(z_{\mathrm{QSO}}+1\right)\times 1216\,\mathrm{\AA}$, ensuring that the Ca II lines fall outside the Lyman alpha forest of the background quasar to avoid contamination. This yields a sample with impact parameter $r_{p}\leq 200\,\mathrm{kpc}$ having 855,703 galaxy-quasar pairs with 623,053 unique BGS galaxies and 496,756 unique QSOs. The median redshift of the unique BGS galaxies is $\langle z\rangle=0.1712$.
Figure 2: Examples of composite spectra as a function of impact parameters. The shades underneath are the associated uncertainties via bootstrapping. The solid blue lines are the best-fit absorption profiles of Ca II K and H. The vertical dashed lines are the central wavelengths of K and H lines. The median impact parameter is displayed in the lower right corner. Note that the scale of y-axis changes as a function of impact parameters.
@@ 2.2 Constructing composite spectra
To extract the absorption lines from the quasar spectra, we first employ a dimensional reduction technique, called nonnegative matrix factorization (NMF) (Lee & Seung, 1999), with code [^3] https://github.com/guangtunbenzhu/NonnegMFPy developed by Zhu & Ménard (2013a) and Zhu (2016). With NMF, we normalize the observed quasar spectrum by the reconstructed continuum, derived from a basis set of NMF eigenspectra built from SDSS DR7 quasars (Schneider et al., 2010; Zhu & Ménard, 2013a), to remove the spectral features intrinsic to the quasar. To further remove fluctuations not captured by the NMF method, we obtain a smooth continuum by applying a median filter with a window size of 71 pixels ($\sim 56\,\rm\AA$) to the NMF normalized spectrum and divide the NMF normalized spectrum by the smooth continuum. The window size is empirically chosen to be large enough to ensure sensitivity to the continuum rather than the absorption itself, which spans about 5 pixels (a full width at half maximum of $\sim\mathrm{3.5\,\AA}$). We then shift the normalized quasar spectra to the rest frame of the foreground galaxies, and resample them onto a common wavelength grid with a resolution matching that of the DESI spectra when the doublet line is redshifted to $z=0.2$ ($\sim\mathrm{0.67\,\AA}$). We obtain composite spectra with a median estimator to mitigate the contributions from spectral outliers that occasionally occur. Finally, to account for a sub-percent of zero-point offset, the final composite spectrum is obtained by normalizing the composite spectrum by the median value of the pixels within 3900 and 4100 $\rm\AA$ with the pixels within 5 $\rm\AA$ of Ca II K and H lines being masked. The uncertainties of the median composite spectra are estimated by bootstrapping the samples 500 times.
@@
Figure 3: Dependence of the Ca II K and H equivalent width on stellar masses as a function of impact parameters. Left: star-forming galaxies. Right: quiescent galaxies. The median stellar masses are listed in the lower left. The blue shaded bands in the right panel correspond to the two highest mass bins for star-forming galaxies in the left panel. The cyan and yellow, square data points are the K line measurements from Rubin et al. (2022) multiplied by a factor of 1.69 to compare with our doublet measurements.

### v2

# A Comprehensive Characterization of Galaxy–cool CGM Connections at $z<0.4$ with DESI Year 1 Data

@@ Abstract
We investigate the relationships between the cool circumgalactic medium (CGM), traced by Ca II absorption lines, and galaxy properties at $z<0.4$ using $\sim 900{,}000$ galaxy–quasar pairs within $200\,\rm kpc$ from the Year 1 data of the Dark Energy Spectroscopic Instrument (DESI). This large data set enables us to obtain composite spectra with sensitivity reaching to the $\rm m\AA$ level and to explore the Ca II absorption as a function of stellar mass, star formation rate (SFR), redshift, and galaxy types, including active galactic nuclei (AGNs). Our results show a positive correlation between the absorption strength and stellar mass of star-forming galaxies with $\langle W_{0}^{\rm Ca\ II}\rangle\propto M_{*}^{0.5}$ over 3 orders of magnitude in stellar mass from $\sim 10^{8}$ to $10^{11}\,M_{\odot}$, while such a mass dependence is weaker for quiescent galaxies. At a fixed mass, Ca II absorption is stronger around star-forming galaxies than quiescent ones especially within impact parameters $30\,\rm kpc$. Among star-forming galaxies, the Ca II absorption further correlates with SFR, following $\propto\mathrm{SFR^{0.3}}$. However, in contrast to the results at higher redshifts, stronger absorption is not preferentially observed along the minor axis of star-forming galaxies, indicating a possible redshift evolution of CGM dynamics resulting from galactic feedback. Moreover, no significant difference between the properties of the cool gas around AGNs and galaxies is detected. Finally, we measure the absorption profiles with respect to the virial radius of dark matter halos and show that the total Ca II mass in the CGM is comparable to the Ca mass in the ISM of galaxies.
@@ 1 Introduction
The circumgalactic medium (CGM) contains signatures of complex physical processes that regulate gas flowing in and out of galaxies and drive galaxy evolution (see Tumlinson et al., 2017; Péroux & Howk, 2020; Faucher-Giguère & Oh, 2023, for reviews on the topic of the CGM). Therefore, observed properties of the CGM and its connection to the properties of galaxies have been considered as crucial constraints on the models of galaxy evolution, especially the subgrid models implemented for the feedback processes driven by the explosions of massive stars and the activities of supermassive black holes (e.g., Ford et al., 2016; Liang et al., 2016; Anglés-Alcázar et al., 2017; Oppenheimer et al., 2018). To probe the properties of the CGM, absorption line spectroscopy has been widely used (e.g., Bahcall & Spitzer, 1969; Bergeron, 1986; Chen et al., 2010; Nielsen et al., 2013; Tumlinson et al., 2013; Lan et al., 2014; Zhu et al., 2014; Anand et al., 2021; Bouché et al., 2025). By observing the spectra of background sources intercepting the CGM of galaxies, one can detect absorption line features induced by various transitions of elements, and investigate the properties of the circumgalactic gas, such as the radial distribution, kinematics, abundance, metallicity, and gas density, and explore how they correlate with the properties of the galaxies (e.g., Zych et al., 2009; Zhu et al., 2014; Lan et al., 2014; Nielsen et al., 2017; Pointon et al., 2019; Lan, 2020; Anand et al., 2021; Lehner et al., 2024; Nateghi et al., 2024; Cherrey et al., 2025).
@@
Multiwavelength observations are required to investigate the galaxy–CGM connections at different redshifts owing to the accessibility of strong absorption-line transitions. At $z<0.3$, UV spectrographs, COS and STIS (Green et al., 2012; Woodgate et al., 1998), on the Hubble Space Telescope have been used to detect absorption species, such as H I, O I, O VI, C II, C IV, Si II, and Si III, tracing gas in various physical conditions (e.g., Tumlinson et al., 2013; Werk et al., 2013, 2014; Bordoloi et al., 2014; Liang & Chen, 2014; Borthakur et al., 2015, 2016; Werk et al., 2016; Prochaska et al., 2017; Chen et al., 2020; Tchernyshyov et al., 2023; Zheng et al., 2024). These measurements have led to new explorations of the nature and origins of the multiphase CGM. However, given the sensitivity of the current instruments, only bright quasars can yield UV spectra with sufficient sensitivity for absorption-line studies (e.g., Tumlinson et al., 2013). This in turn limits the number of galaxy–quasar pairs and the parameter space of galaxy properties available for galaxy–CGM correlation analysis. On the other hand, at $z\gtrsim 0.4$, optical spectra can be used to investigate the connection between galaxies and the cool CGM traced by Mg II absorption lines, one of the key accessible absorption species at optical wavelengths (e.g., Nestor et al., 2005; Chen et al., 2010; Nielsen et al., 2013; Zhu & Ménard, 2013a; Anand et al., 2021; Huang et al., 2021; Chen et al., 2023; Guha et al., 2024; Zou et al., 2024; Bouché et al., 2025). With large optical spectroscopic data sets provided by sky surveys, such as the Sloan Digital Sky Survey (SDSS; York et al. 2000) and the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2013), statistical samples, consisting of thousands of galaxy–quasar pairs, have been built and used to measure the properties of the cool CGM as a function of galaxy properties, including stellar mass, star formation rate (SFR) and active galactic nucleus (AGN) activity, across redshifts (e.g., Zhu & Ménard, 2013a; Lan et al., 2014; Lan, 2020; Anand et al., 2021; Chang et al., 2024; Wu et al., 2024), revealing correlations between the cool CGM and galaxy properties. However, observational measurements of these correlations at low redshifts are still limited, and therefore whether the relationships between galaxies and the cool CGM detected at $z\gtrsim 0.4$ exist at lower redshifts remains to be addressed.
@@
[…]
@@ 2.1 DESI galaxy and quasar catalogs
The DESI project is designed to measure the expansion rate of the Universe across cosmic time by using galaxies, quasars, and gas as tracers of large-scale structure at different redshifts (Levi et al., 2013; DESI Collaboration et al., 2022). To this end, a dedicated spectroscopic observational program has been developed utilizing an instrument consisting of 10 spectrographs and 5000 fibers (Silber et al., 2023; Miller et al., 2024). DESI primarily obtains spectra of four types of extragalactic sources at different redshift ranges, including galaxies that can be observed during the bright time targeted by the Bright Galaxy Survey ($z<0.6$; Hahn et al. 2023), luminous red galaxies ($0.4<z<1.0$; Zhou et al. 2023), emission-line galaxies ($0.6<z<1.6$; Raichoor et al. 2023) and quasars ($z<3$; Chaussidon et al. 2023). These sources are selected based on the images of the DESI Legacy Imaging Surveys (Dey et al., 2019). The wavelength coverage of DESI spectra is from 3600 to 9800 Å with the spectral resolution varying from $R\sim 2000$ to 5000 for the blue camera to the NIR camera. The raw spectroscopic data are processed with a pipeline developed by Guy et al. (2023). In addition, an automatic pipeline, called redrock [^1] https://github.com/desihub/redrock, is developed and used for determining the redshifts of the sources from the spectra (Bailey et al. 2024; see also Anand et al. 2024). In order to maximize the efficiency of the survey, dedicated algorithms and codes have also been developed for target selection (Myers et al., 2023), fiber assignments (Raichoor et al., 2024) and survey operations (Schlafly et al., 2023). Prior to the main spectroscopy survey, DESI has undergone the “Survey Validation (SV)” (DESI Collaboration et al., 2024a) to validate the performance of the target selections. The five-month spectroscopic data from Survey Validation was released in 2023 June as part of Early Data Release (DESI Collaboration et al., 2024b).
@@
Figure 1: SFR–mass distribution of the foreground galaxies. The solid line divides the samples into star-forming and quiescent galaxies. The contours mark the levels of numbers at 10, 100, 500, 1000, 5000, and 8000.
@@
In this work, we use the Year 1 data set of the DESI survey (DESI Collaboration et al., 2025a), which has yielded new measurements for constraining the expansion history of the Universe from $z=0$ to 2 (DESI Collaboration et al., 2024c, d, e, f, g, h, 2025b, 2025c). More specifically, we construct our foreground galaxy and background source samples with the following two components:
@@
  • Bright Galaxy Survey (BGS): We use galaxies with $0.02<z_{\mathrm{BGS}}<0.4$ observed in the bright galaxy survey as our foreground galaxies. The BGS sample consists of two selections, BGS bright and BGS faint samples. The BGS bright sample is selected with $r<19.5$ mag, and the BGS faint sample is selected with magnitude and color cuts (Hahn et al., 2023). Moreover, following the criteria from Hahn et al. (2023) and Lan et al. (2023), we select galaxies with $\Delta\chi^{2}\geq 40$ (the difference of $\chi^{2}$ between the second and the first best-fit spectral models from redrock).
@@
    Besides the redshift information, we use the physical properties of the BGS galaxies, including the stellar mass ($M_{*}$) and SFR, provided by Siudek et al. (2024). The authors use the CIGALE package (Boquien et al., 2019) for spectral energy distribution fits to observed magnitudes in the $g$, $r$, $z$, WISE 1 and WISE 2 bands extracted by the Tractor algorithm (Lang et al., 2016). The final fitting results exclude sources with low signal-to-noise ratio in photometry and poor fitting quality. For details, we refer the readers to Siudek et al. (2024).
@@
    As shown in the figure, the BGS sample includes star-forming galaxies with $M_{*}$ down to $10^{7}-10^{8}\,M_{\odot}$ allowing us to probe the CGM of galaxies with mass across 3–4 orders of magnitude. The median mass uncertainty of the sample is 0.12 dex. For SFR, the typical uncertainties are 0.3 dex for star-forming galaxies and 8.1 dex for quiescent galaxies. We note that the reported uncertainty is in logarithmic scale. For quiescent galaxies with low estimated SFR ($\log_{10}\mathrm{SFR}/M_{\odot}\rm yr^{-1}\sim-7$), the corresponding uncertainty in linear scale is $\sim 10^{-6}M_{\odot}\rm yr^{-1}$. This precision is sufficient for classifying galaxies into star-forming and quiescent ones.
    In addition, we construct an AGN sample by selecting sources located in the AGN and composite regions (Kewley et al., 2006) of the BPT diagram (Baldwin et al., 1981). For this selection, we require a signal-to-noise ratio greater than 2 for the $\rm H\alpha$, $\rm H\beta$, [O III] 5007, and [N II] 6584 emission lines. The flux measurements for these lines are provided by the FastSpecFit [^2] https://fastspecfit.readthedocs.io/en/latest/ Spectral Synthesis and Emission-Line Catalog (FastSpecFit 2.1; Moustakas et al. 2023).
  • Quasars: For the background sources, we use the quasar sample that is selected with a random forest algorithm combining magnitude and color information (Chaussidon et al., 2023). Similar to the BGS sample selection, we apply $\Delta\chi^{2}\geq 20$ to obtain sources with robust redrock redshift measurements (Alexander et al., 2023).
[…]
Figure 2: Examples of composite spectra as a function of impact parameters. The shades underneath are the associated uncertainties via bootstrapping. The solid blue lines are the best-fit absorption profiles of Ca II K and H. The vertical dashed lines are the central wavelengths of K and H lines. The median impact parameter is displayed in the lower right corner. Note that the scale of the y-axis changes as a function of impact parameters.
@@ 2.2 Constructing composite spectra
[…]
@@
Figure 3: Dependence of the Ca II K and H equivalent width on stellar masses as a function of impact parameters. Left: star-forming galaxies. Right: quiescent galaxies. The median stellar masses are listed in the lower left. The blue shaded bands in the right panel correspond to the two highest mass bins for star-forming galaxies in the left panel. The cyan and yellow square data points are the K line measurements from Rubin et al. (2022) multiplied by a factor of 1.69 to compare with our doublet measurements.